OpenStudy (anonymous):

GloryRae: graph of y = |x – 1| – 1

OpenStudy (anonymous):

vertex at 1, -1 line of symmetry at x = 1

OpenStudy (sriram):

x>1 y=x-2 x<1 y=-x

OpenStudy (anonymous):

x = 1 x<1 and y=-x I Believe! (:

OpenStudy (sriram):



OpenStudy (sriram):

the point where they meet=1,-1